# From Prejudice to Parity: A New Approach to Debiasing Large Language Model Word Embeddings

### v4

@@ 6.1 Mean Average Cosine Similarity
In our research, we have chosen to utilize Mean Average Cosine Similarity (MAC) as our primary metric for assessing bias in word embeddings. This decision is informed by the work of Manzini et al. (2019), who posit that MAC can be viewed as an extension of the Word Embedding Association Test (WEAT), specifically adapted for a multiclass setting. The MAC and WEAT serve distinct, yet complementary purposes. While WEAT is designed to focus on specific associations between word vectors and predefined concepts (such as gender or race), MAC provides a broader perspective by measuring overall similarity patterns across different groups. This makes MAC less sensitive to specific word choices, thereby revealing biases that might be overlooked by WEAT. In essence, both metrics contribute to a comprehensive understanding of bias in word embeddings. However, the use of MAC is particularly beneficial in our research as it complements the findings of WEAT, providing a more holistic view of bias in the data. This approach allows us to capture a wider range of biases, thereby enhancing the robustness of our analysis.
@@ 6.2 Stereotype Score
Our research focuses on evaluating and mitigating stereotypical bias in NLI tasks using the Stereoset dataset. This dataset comprises pairs of sentences differing only in the substitution of words related to social groups like gender, race, or religion. The objective is to predict their relationship as same, entailment, or contradiction. We introduce a method aimed at reducing bias in word embeddings, with Stereotype Score $SS$ values closer to 50 indicating decreased bias. Table 2 presents DeepSoftDebias’s results alongside existing approaches on the Stereoset dataset. Notably, DeepSoftDebias achieves the lowest $SS$ across all social groups, demonstrating its effectiveness in bias reduction. Particularly impressive is DeepSoftDebias’s performance in the gender and race categories, where it significantly outperforms existing methods. For instance, with the SFR-Embedding-Mistral Jiang et al. (2023) model, DeepSoftDebias achieves an $SS$ of 50 for gender and 50.409 for race using the Llama-2-7b model. Additionally, DeepSoftDebias attains a score of 51.282 for the Zephyr-7b-beta Tunstall et al. (2023) or 48.717 for Alpaca-7b Taori et al. (2023). We present these score in 2 an illustration of these scores in Fig. 6.
@@ 6.3 Crows-Pairs Dataset
Our study evaluates social bias in natural language generation tasks using the CrowS Pairs dataset, comprising pairs of sentences differing in their degree of bias. By ranking these sentences according to bias level, we quantify the effectiveness of various methods in reducing bias in word embeddings. But as our work is based on word embeddings instead of getting the log-likelihood of the next token from the language model, we compute the average sentence vector for the common parts shared between two sentences. Next, we compare the similarity of this average sentence vector with the uncommon part (i.e., the modified tokens) using word embeddings. By doing so, we capture the semantic differences between stereotypical and non-stereotypical components within the sentence pairs. The rest of the metric remains the same.
@@
where Crows Pair Stereotype Score (CSS) is the number of stereotypical samples that agree with their label direction and Crows Pairs Anti-stereotype Score (CAS) is the number of anti-stereotypical samples that agree with their label direction. Label direction refers to the label given the pair of sentences whether they are stereotypical or anti-stereotypical. In our evaluation we get the average sentence vector of the context and the more and less (anti-)stereotypical sentence. We then see whether the context vector is closer to the more (anti-)stereotypical sentence or the less (anti-)stereotypical sentence. If it is closer to the more (anti–)stereotypical sentence,, then we state that it agrees with the (anti–)stereotype, i.e., the label direction. Particularly noteworthy is DeepSoftDebias’s superior performance in the gender and religion categories. For instance, with the Yi-6B model, DeepSoftDebias achieves a score of 49.62 for gender and 50.48 for religion with the falcon-7b model. Similarly, using the SFR-Embedding-Mistral model, DeepSoftDebias achieves a score of 49.03 for race biasing the SFR-Embedding-Mistral model. These results underscore the effectiveness of DeepSoftDebias in mitigating social bias in word embeddings. We present these score in 3 and depict the variation of these scores in Fig. 8.
We also report the CSS and CAS score which refer to the CrossNER Stereotype score, i.e., the number of times the model agrees with the more stereotypes statement when the label direction is stereotype, and the CrossNER Anti-stereotype score, which refers to the number of times the model agrees with the more anti-stereotyped statement when the label direction was anti-sterotype.
@@ 7 Discussion
  • A single layer neural network can effectively de-bias embeddings with $\text{dim}\leq 1024$.
  • A two-layer neural network can effectively debias embeddings with $\text{dim}\leq 2048$.
  • A two-layer neural network with an increased layer size can effectively de-bias embeddings with $\text{dim}\leq 4096$.
  • A three-layer neural network can effectively debias embeddings with $\text{dim}\leq 4450$.
As a step for future work, we are curious to investigate scaling patterns to a further extent.A visualization of this is provided in Fig 11
@@
We find that while DeepSoftDebias is effective at reducing bias across gender, race, and religion. We conclude this after testing on multiple embeddings, and multiple datasets and evaluating on multiple performance metrics. As a step for future work, we are curious to investigate whether our proposed approach works towards other forms of bias as well.
@@ 9 Limitations
While we do perform exhaustive analysis to test our proposed methodology, our study is monolingual and covers datasets only in English. Consequently, our downstream tasks are also tested only in English. Further, we were unable to conduct test on API-based models at this time. Our testing was also constrained by the limitations of GPU VRAM, which prevented us from extending our testing to larger models such as Llama-65B. These models could not be accommodated within the GPU VRAM, even after applying quantization to 8 bits. Consequently, the largest model that we were able to test was the Gemma-7B model.
@@ Appendix A Table of words and bias they represent
Asian "manager", "executive", "redneck", "hillbilly", "leader", "farmer"
Caucasian "doctor", "engineer", "laborer", "teacher"
@@ Appendix D Downstream Testing Results
Model Variant Baseline Debiased Text Class. Acc. DSB Debiased Text Class. Acc. Baseline Debiased NER Macro F1 Avg. DSB Debiased NER Macro F1 Avg.
Gemma gemma-2b 0.7655 0.7964 0.469 0.484
BAAI bge-base-en-v1.5 0.8296 0.822 0.458 0.421
Mistral SFR-Embedding-Mistral 0.8297 0.821 0.404 0.428
Gemma gemma-7b 0.516 0.8032 0.198 0.475
Zephyr zephyr-7b-beta 0.7997 0.81 0.403 0.429
mxbai mxbai-embed-large-v1 0.8366 0.7903 0.461 0.455
Table 5: Downstream testing results comparison with embeddings debiased using DeepSoftDebias and the baseline SoftDebais Method. The first two columns represent results for downstream performance on sentiment analysis. The second two columns represent results for downstream performance on NER.
@@ D.1 Sentiment Classification
In our study, we employ downstream testing to assess the utility of embeddings debiased using DeepSoftDebias across two key natural language processing tasks: text classification and named entity recognition (NER). Utilizing the IMDB Sentiment Classification dataset İlhan Tarımer et al. (2019) and Stanford TreeBank Dataset for text classification, featuring labeled movie reviews as positive or negative, we compute the average sentence vectors using both original and debiased embeddings. Training XGBoost Chen and Guestrin (2016) classifiers on these vectors, we compare their accuracy on the test set, recognizing accuracy as a straightforward metric for binary classification tasks like sentiment analysis. Notably, our results reveal a performance improvement when debiasing in the gender and religion directions, whereas a slight decrease in performance is observed in the case of race debiasing. We provide these results in Table:6 for IMDB Sentiment classification and Table:7 for Stanford Sentiment Treebank. A visual representation of these results in Fig. 9.
D.2 Named Entity Recognition (NER)
In our research, we examine the performance of debiased embeddings in the domain of named entity recognition (NER) using the Reuters subset of the CrossNER Liu et al. (2020) dataset. This dataset comprises news domain sentences annotated with four entity types: person, location, organization, and product. Employing a simple BiLSTM model, we input padded arrays of embeddings for each sentence and trained the model on the dataset. We evaluate the models’ performance on the test set using the macro-averaged F1-score, a metric that balances precision and recall, crucial for accurate entity identification and classification. To mitigate potential bias towards more frequent entity types, we adopt macro-averaging, allotting equal importance to each entity type. Remarkably, our findings indicate a slight performance boost when using debiased embeddings in all three directions compared to biased embeddings. We provide these results in Table:6and a visual representation of these results in Fig. 10.
D.3 Semantic Textual Similarity

### v5

@@ 6.1 Mean Average Cosine Similarity
In this work, we chose to utilize Mean Average Cosine Similarity (MAC) as our primary metric for assessing bias in word embeddings. This decision is informed by the work of Manzini et al. (2019), who posit that MAC can be viewed as an extension of the Word Embedding Association Test (WEAT), specifically adapted for a multiclass setting. While WEAT is designed to focus on specific associations between word vectors and predefined concepts (such as gender or race), MAC provides a broader perspective by measuring overall similarity patterns across different groups. This makes MAC less sensitive to specific word choices, thereby revealing biases that might be overlooked by WEAT. In essence, both metrics contribute to a comprehensive understanding of bias in word embeddings. However, the use of MAC is particularly beneficial in our research as it complements the findings of WEAT, providing a more holistic view of bias in the data. This approach allows us to capture a wider range of biases, thereby enhancing the robustness of our analysis.
@@ 6.2 Stereotype Score
Our research focuses on evaluating and mitigating stereotypical bias in NLI tasks using the Stereoset dataset. This dataset comprises pairs of sentences differing only in the substitution of words related to social groups like gender, race, or religion. The objective is to predict their relationship as same, entailment, or contradiction. We introduce a method aimed at reducing bias in word embeddings, with Stereotype Score $SS$ values closer to 50 indicating decreased bias. Table 2 presents DeepSoftDebias’s results alongside existing approaches on the Stereoset dataset. Notably, DeepSoftDebias achieves the lowest $SS$ across most social groups, demonstrating its effectiveness in bias reduction. Particularly impressive is DeepSoftDebias’s performance in the gender and race categories, where it significantly outperforms existing methods. For instance, with the Mamba 2.8b Jiang et al. (2023) model, DeepSoftDebias achieves an $SS$ of 50 for gender and with the Llama 3 8b model it acheives a SS of 49.8 for race. We present these scores in Table 2 and an illustration of these scores in Fig. 6.
@@ 6.3 Crows-Pairs Dataset
Our study also evaluates social bias in natural language generation tasks using the CrowS Pairs dataset, comprising pairs of sentences differing in their degree of bias. By ranking these sentences according to bias level, we quantify the effectiveness of various methods in reducing bias in word embeddings. But as our work is based on word embeddings instead of getting the log-likelihood of the next token from the language model, we compute the average sentence vector for the common parts shared between two sentences. Next, we compare the similarity of this average sentence vector with the uncommon part (i.e., the modified tokens) using word embeddings. By doing so, we capture the semantic differences between stereotypical and non-stereotypical components within the sentence pairs. The rest of the metric remains the same.
@@
where Crows Pair Stereotype Score (CSS) is the number of stereotypical samples that agree with their label direction and Crows Pairs Anti-stereotype Score (CAS) is the number of anti-stereotypical samples that agree with their label direction. Label direction refers to the label given the pair of sentences whether they are stereotypical or anti-stereotypical. In our evaluation we get the average sentence vector of the context and the more and less (anti-)stereotypical sentence. We then see whether the context vector is closer to the more (anti-)stereotypical sentence or the less (anti-)stereotypical sentence. If it is closer to the more (anti-)stereotypical sentence2901, then we state that it agrees with the (anti-)stereotype, i.e., the label direction. Particularly noteworthy is DeepSoftDebias’s superior performance in the gender and religion categories. For instance, with the mamba-1.4b model, DeepSoftDebias achieves a score of 50.38 for gender and 50.48 for religion with the mamba-2.8b model. Similarly, using the bge-base-en v1.5 model, DeepSoftDebias achieves a score of 50.19 for debiasing for the bias-type of race. These results underscore the effectiveness of DeepSoftDebias in mitigating social bias in word embeddings. We present these scores in 3 and 5 and depict the variation of these scores in Fig. 8.
@@
6.4 Downstream Testing
For our downstream evaluation, we utilized the GLUE benchmark Wang et al. (2018). We present the performance differentials between the original word embeddings and their debiased counterparts, processed through various debiasing methods focusing on the three categories of GLUE tasks: single-sentence tasks, sentence-pair tasks, and inference-based tasks. For the single-sentence task, specifically the Stanford Sentiment Treebank (SST), we report accuracy as the primary metric in Table 6. In the case of sentence-pair tasks, exemplified by the Microsoft Research Paraphrase Corpus (MRPC), we use the F1 score as our performance indicator in Table 8. For the inference-based tasks, which include QNLI (Question-answering Natural Language Inference), WNLI (Winograd Natural Language Inference), RTE (Recognizing Textual Entailment), and MNLI (Multi-Genre Natural Language Inference), we report the average delta of the F1 scores across these four tasks in Table 7. For inference-based tasks we see a average gain in performance of 0.017 F1 Score for gender, followed by 0.027 for religion, and a net average performance delta of 0 for race. For sentence pair tasks, we see a performance delta of 0.047 for gender, 0.035 for race and -0.037 F1 score for religion. Our results indicate that the DeepSoft Debias method yields performance comparable to that of the original biased word embeddings, with variations in performance ranging from approximately 2% to 4% across the different metrics. We hypothesize that this outcome is attributable to the varying degrees of bias present in the datasets, which the models have become less susceptible to following the debiasing process. We have provided a detailed score $\Delta$s for all the models we have applied out DeepSoftDebias debiasing method on with the GLUE Tasks in Tables 9, 10, 11, 12, 13, 14 and 16 in the Appendix.
@@ 7 Discussion
  • A single residual block(RB) neural network can effectively de-bias embeddings with $\text{dim}\leq 1024$.
  • A two RB neural network can effectively debias embeddings with $\text{dim}\leq 2048$.
  • A three RB neural network with an increased layer size can effectively de-bias embeddings with $\text{dim}\leq 4450$.
With respect to future work, we are curious to investigate scaling patterns to a further extent. A visualization of this is provided in Fig 11. We offer a further detailed discussion on hyperparameter tuning the DeepSoftDebias Method in Question 5 of our FAQ section. 2
@@
We find that while DeepSoftDebias is effective at reducing bias across gender, race, and religion. We conclude this after testing on multiple embeddings, and multiple datasets and evaluating on multiple performance metrics. As a step for future work, we are curious to investigate whether our proposed approach works towards other forms of bias as well. We offer a further detailed quantitative discussion of the bias types it works best with, in Question 6 of our FAQ Section.
@@ 9 Limitations
While we do perform exhaustive analysis to test our proposed methodology, our study is monolingual and covers datasets only in English. Consequently, our downstream tasks are also tested only in English. Further, we were unable to conduct test on API-based models at this time. Our testing was also constrained by the limitations of GPU VRAM, which prevented us from extending our testing to larger models such as Llama-65B. These models could not be accommodated within the GPU VRAM, even after applying quantization to 8 bits. Consequently, the largest model that we were able to test was the Gemma-2-9b model.
@@ Appendix A Table of words and bias they represent
Caucasian "manager", "executive", "redneck", "hillbilly", "leader", "farmer"
Asian "doctor", "engineer", "laborer", "teacher"
@@ Appendix C Stereoset Scores of DeepSoftDebias
We present the MAC scores, Stereotype Scores, Crows-Pairs Metric Scores in Table 5
Bias Type Model Name Biased MAC Soft-Debiased MAC DSD MAC Soft-Debiased SS DSD SS Soft-Debiased CMS DSD CMS
gender gte-Qwen2-7B-instruct 0.469 - 0.95 25.21 45.04 32.82 41.22
bge-base-en-v1.5 0.447 0.88 0.966 47.93 43.39 43.13 45.04
bge-large-en-v1.5 0.408 0.881 0.921 49.59 48.35 41.98 50.76
roberta-base 0.025 0.179 0.922 52.06 50.41 53.44 51.53
bert-base-uncased 0.243 0.408 0.921 50.41 52.07 40.84 45.8
bert-large-uncased 0.324 0.734 0.917 52.89 48.76 50 48.85
gemma-2-2b 0.46 - 0.894 25.62 49.59 26.72 37.4
gemma-2-9b 0.502 - 0.915 25.21 48.76 30.15 46.95
gemma-2b 0.056 - 0.847 19.42 47.11 28.24 41.98
gemma-7b 0.54 0.16 0.908 45.45 48.35 45.04 45.8
GritLM-7B 0.38 - 0.905 24.38 49.59 25.95 39.31
Meta-Llama-3-8B 0.254 - 0.939 19.83 50 23.66 49.24
gpt2-xl 0.497 0.91 0.945 49.17 48.35 46.18 49.62
SFR-Embedding-2_R 0.388 - 0.906 26.03 51.24 30.53 40.08
mamba-1.4b-hf 0.342 - 0.935 43.39 46.28 40.08 50.38
mamba-2.8b-hf 0.101 - 0.919 21.9 50 18.32 51.91
race gte-Qwen2-7B-instruct 0.426 - 0.971 26.02 51.23 16.86 63.18
bge-base-en-v1.5 0.467 0.903 0.987 51.84 51.02 60.27 50.19
bge-large-en-v1.5 0.424 0.938 0.99 51.33 51.43 39.34 50.78
roberta-base 0.025 0.205 0.863 48.77 47.44 30.62 47.04
bert-base-uncased 0.245 0.448 0.974 52.25 51.43 29.07 46.07
bert-large-uncased 0.354 0.798 0.997 49.28 49.69 61.43 44.96
gemma-2-2b 0.44 - 0.966 29.3 47.95 41.67 56.2
gemma-2b 0.046 - 0.838 21.31 50.31 12.98 51.16
gemma-7b 0.546 0.111 0.951 47.75 46.72 53.1 43.22
GritLM-7B 0.417 - 0.971 21.41 51.23 15.5 46.12
Meta-Llama-3-8B 0.252 - 0.937 22.23 49.8 27.91 43.99
gpt2 0.01 - 0.822 20.49 48.77 12.79 38.18
gpt2-xl 0.477 0.927 0.977 49.59 51.02 50.39 55.62
SFR-Embedding-2_R 0.44 - 0.99 29.2 51.74 32.75 59.88
mamba-1.4b-hf 0.356 - 0.994 45.7 49.39 56.98 47.29
mamba-2.8b-hf 0.109 - 0.958 21.11 51.33 19.19 52.52
religion gte-Qwen2-7B-instruct 0.442 - 0.975 21.79 57.69 19.05 70.48
bge-base-en-v1.5 0.474 0.897 0.987 50 47.44 49.52 61.9
bge-large-en-v1.5 0.412 0.895 0.94 53.85 55.13 60 56.19
roberta-base 0.025 0.209 0.936 51.28 56.41 36.19 34.29
bert-base-uncased 0.276 0.477 0.991 56.41 46.15 32.38 68.57
bert-large-uncased 0.312 0.745 0.971 52.56 48.72 42.86 60.95
gemma-2-2b 0.523 - 0.957 20.51 52.56 29.52 44.76
gemma-2b 0.095 - 0.933 24.36 55.13 12.38 58.1
gemma-7b 0.372 - 0.907 43.59 52.56 40.95 55.48
GritLM-7B 0.485 - 0.963 17.95 56.41 25.71 55.24
Meta-Llama-3-8B 0.264 - 0.951 20.51 58.97 20 52.43
gpt2 0.018 - 0.953 12.82 52.56 22.86 63.81
SFR-Embedding-2_R 0.455 - 0.959 34.62 55.13 41.9 54.29
mamba-1.4b-hf 0.267 - 0.97 55.13 55.13 41.9 49.52
mamba-2.8b-hf 0.087 - 0.945 19.23 46.15 27.62 50.48
Table 5: Debiasing Results on LLMs: MAC, CMS and SS on all the models we have tested our method on
@@ D.1 Text Classification
In our study, we employ downstream testing to assess the utility of embeddings debiased using DeepSoftDebias across two key natural language processing tasks: text classification and named entity recognition (NER). Utilizing the Stanford TreeBank DatasetSocher et al. (2013) for text classification. Training XGBoost Chen and Guestrin (2016) classifiers on these vectors, we compare their accuracy on the test set, recognizing accuracy as a straightforward metric for binary classification tasks like sentiment analysis. Notably, our results reveal a slight performance improvement when debiasing in the gender and religion directions inmost cases, whereas a slight decrease in performance is observed in the case of race debiasing in mostr cases. We provide these results in Table:9 for Stanford Sentiment Treebank. A visual representation of these results in Fig. 9.
D.2 Semantic Textual Similarity